OpenStudy (anonymous):

Solve for x in the proportion:

OpenStudy (anonymous):

\[\frac{ x }{ x + 12 }\] = \[\frac{ 1 }{ x - 1 }\]

OpenStudy (anonymous):

@Luigi0210

OpenStudy (johnweldon1993):

So you would first cross multiply

OpenStudy (luigi0210):

@johnweldon1993 's got you brah

OpenStudy (anonymous):

x + 12 = 2x - 1

OpenStudy (johnweldon1993):

Not quite....remember x * x = x² try that 1 more time

OpenStudy (anonymous):

x + 12 = x^2 - 1 :P

OpenStudy (luigi0210):

Try again >.>

OpenStudy (johnweldon1993):

Closer x(x -1) x * x = ? x * -1 = ?

OpenStudy (anonymous):

x + 12 = x^2 - x?

OpenStudy (johnweldon1993):

There we go... Okay ...next step is to combine like terms

OpenStudy (anonymous):

cancel out x on each side?o;

OpenStudy (johnweldon1993):

Right....lets subtract an 'x' from each side...what do we get when we do that?

OpenStudy (anonymous):

12 = x^2 - 2x?

OpenStudy (johnweldon1993):

Right....now that would be annoying to solve....so now lets subtract 12 from both sides... x² - 2x - 12 = 0 Look enough like a quadratic?

OpenStudy (anonymous):

Yes.

OpenStudy (johnweldon1993):

So now lets use the quadratic formula to solve... \[\large \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] Using a = 1 b = -2 c = -12 Can you do that?

OpenStudy (anonymous):

Let me plug it in. gimme a sec.

OpenStudy (johnweldon1993):

Take your time :P

OpenStudy (anonymous):

\[ \frac{ 2 \pm \sqrt{52}}{ 2 }\]

OpenStudy (johnweldon1993):

Perfect so now you will have 2 answers *notice the +/- sign 2 equations you need to solve are \[\large \frac{ 2 + \sqrt{52} }{ 2 }\] and \[\large \frac{ 2 - \sqrt{52} }{ 2 }\]

OpenStudy (anonymous):

9.21 / 2 -7.21/2

OpenStudy (anonymous):

\[\approx \]

OpenStudy (anonymous):

ops i meant -5.21 :p

OpenStudy (johnweldon1993):

Ahh there you go...lol I was going to say hmm not quite.... but yes 9.21 / 2 and -5.21 / 2 These will be your *approximate* answers for 'x'

OpenStudy (anonymous):

I got it (: Thank you!

OpenStudy (johnweldon1993):

No problem!